Find the center of mass of four 2kg bodies placed at eachcorner of a square



of side 1.5 m.

1
Expert's answer
2021-11-07T19:26:36-0500

Let the left bottom body be at the origin.



"4mX=m(0+0+1.5+1.5)\\X=0.75\ m
\\4mY=m(0+0+1.5+1.5)\\Y=0.75\ m\\"

COM is at center of square.